# Unlocking hidden biomolecular conformational landscapes in diffusion models at inference time

### v1

@@ 4.4 Rapid free energy estimation with BioEmu
To illustrate the ability of ConforMix to improve sampling in a variety of situations, we implement it in BioEmu [lewis_scalable_2025] as well. Unlike Boltz, BioEmu produces broad distributions of conformations with default sampling. However, default sampling is an inefficient way to observe rare events or estimate quantities such as free energy differences, so we assess the ability of ConforMix to rapidly estimate free energies from BioEmu.
@@ 6 Acknowledgments
Unlocking hidden biomolecular conformational landscapes in diffusion models at inference time
Daniel D. RichmanThese authors contributed equally to this work.
Stanford University
ddrichma@stanford.edu &Jessica Karaguesian11footnotemark: 1
Stanford University
jkara@stanford.edu Carl-Mikael Suomivuori
Stanford University
carl-mikael.suomivuori@yale.edu &Ron O. Dror
Stanford University
rondror@cs.stanford.edu
@@ Appendix A Datasets
  • Domain motion: 38 proteins consisting of the non-overlapping set of 22 domain motion proteins curated by [lewis_scalable_2025] and 23 open-closed (OC23) conformation proteins curated by [kalakoti_afsample2_2025].
  • Membrane transporters: 15 proteins from the transporter protein set (TP16) curated by [kalakoti_afsample2_2025]. One protein, SPF1, was excluded for compute considerations due to its long sequenve length.
  • Cryptic pockets: 33 proteins that undergo cryptic pocket formation curated by [lewis_scalable_2025]. RMSD metrics were computed using the cryptic pocket region defined by [lewis_scalable_2025] . One protein, Q16539, was excluded because the region definitions were not available.
  • Fold switching: 15 proteins that exhibit fold switching were randomly selected from the set of proteins curated by [porter_extant_2018] and used to benchmark other conformational prediction methods including CF-random ([lee_large-scale_2025]).
@@ Appendix E Outliers
Outlier analysis was performed using default settings of the ramalyze module of the cctbx toolkit, associated with Phenix ([liebschner_macromolecular_2019]), to identify Ramachandran outliers. We also evaluate backbone bond length and bond angle outliers as deviations more than $4\sigma$ from reference values. We assess every generated frame for outliers and compute the total fraction of bonds/angles/dihedrals that are marked as outliers, across all frames and proteins (Table S3). While the fraction of outliers is somewhat elevated in ConforMix sampling, the vast majority of bond lengths, bonds angles, and Ramachandran angles are within physical parameters. ConforMix-generated samples are generally usable as is for visual examination or motion statistics. If needed, most outliers are easily resolvable with a brief relaxation, as is standard in the AlphaFold workflow. For ConforMix, we also report results after our filtering, since we intentionally push some proteins beyond their normal range of flexibility in order to capture the full range of motion and reject unphysical samples afterwards.
@@ Appendix F Twisted diffusion sampling algorithm
The subroutine of twisted diffusion sampling is described below. The Weight function accounts for the effects of the bias potential, and Resample outputs a new subset of the particles based on the weights. We use systematic resampling in our implementation. A more complete description of the algorithm is provided in [wu_practical_2024].
@@ Appendix H Supplemental Data and Figures
Worst-matched
reference
conformation
(harder task)
Default Boltz sampling 0.37 (±0.09) 0.16 (±0.11) 0.23 (±0.08) 0.13 (±0.10)
@@
Best-matched
reference
conformation
(easier task)
Default Boltz Sampling 0.77 (±0.04) 0.49 (±0.14) 0.80 (±0.05) 0.54 (±0.19)
@@
Figure S5: Fill ratio analysis across methods and datasets. The fill ratio, developed in [kalakoti_afsample2_2025], is a TM-score–based metric that quantifies the amount of structure diversity between reference states that is captured.

### v2

@@ 4.4 Rapid free energy estimation with BioEmu
To illustrate the ability of ConforMix to improve sampling in a variety of situations, we implement it in BioEmu [17] as well. Unlike Boltz, BioEmu produces broad distributions of conformations with default sampling. However, default sampling is an inefficient way to observe rare events or estimate quantities such as free energy differences, so we assess the ability of ConforMix to rapidly estimate free energies from BioEmu.
@@ Appendix A Datasets
  • Domain motion: 38 proteins consisting of the non-overlapping set of $2$ domain motion proteins curated by [17] and 23 open-closed (OC23) conformation proteins curated by [13].
  • Membrane transporters: 15 proteins from the transporter protein set (TP16) curated by [13]. One protein, SPF1, was excluded for compute considerations due to its long sequenve length.
  • Cryptic pockets: 33 proteins that undergo cryptic pocket formation curated by [17]. RMSD metrics were computed using the cryptic pocket region defined by [17] . One protein, Q16539, was excluded because the region definitions were not available.
  • Fold switching: 15 proteins that exhibit fold switching were randomly selected from the set of proteins curated by [27] and used to benchmark other conformational prediction methods including CF-random ([15]).
@@ Appendix E Outliers
Outlier analysis was performed using default settings of the ramalyze module of the cctbx toolkit, associated with Phenix ([18]), to identify Ramachandran outliers. We also evaluate backbone bond length and bond angle outliers as deviations more than $4\sigma$ from reference values. We assess every generated frame for outliers and compute the total fraction of bonds/angles/dihedrals that are marked as outliers, across all frames and proteins (Table S3). While the fraction of outliers is somewhat elevated in ConforMix sampling, the vast majority of bond lengths, bonds angles, and Ramachandran angles are within physical parameters. ConforMix-generated samples are generally usable as is for visual examination or motion statistics. If needed, most outliers are easily resolvable with a brief relaxation, as is standard in the AlphaFold workflow. For ConforMix, we also report results after our filtering, since we intentionally push some proteins beyond their normal range of flexibility in order to capture the full range of motion and reject unphysical samples afterwards.
@@ Appendix F Twisted diffusion sampling algorithm
The subroutine of twisted diffusion sampling is described below. The Weight function accounts for the effects of the bias potential, and Resample outputs a new subset of the particles based on the weights. We use systematic resampling in our implementation. A more complete description of the algorithm is provided in [38].
@@ Appendix H Supplemental Data and Figures
Worst-matched reference conformation (harder task) Default Boltz sampling 0.37 (±0.09) 0.16 (±0.11) 0.23 (±0.08) 0.13 (±0.10)
@@
Best-matched reference conformation (easier task) Default Boltz Sampling 0.77 (±0.04) 0.49 (±0.14) 0.80 (±0.05) 0.54 (±0.19)
@@
Figure S5: Fill ratio analysis across methods and datasets. The fill ratio, developed in [13], is a TM-score–based metric that quantifies the amount of structure diversity between reference states that is captured.